OpenStudy (anonymous):

PLEASE HELP. Find the distance between points M(6,16) and Z(-1,14) to the nearest tenth. Can you please show steps? thanks! it deleted the help i got?

OpenStudy (anonymous):

\[\sqrt{(16-14)^2+(6--1)^2}=\sqrt{53}=7.28011 \]

OpenStudy (mrnood):

 you can make right angle triangle as shown work out the vertiacla and horizontal sides (using the points) and then use Pythagoras to work out the length

OpenStudy (mrnood):

OR just copy someone else's answer that they gave above

OpenStudy (anonymous):

Thanks!